40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Jun 10 15:54:14 2015, Steve, Update, VAC, Vacuum comp. rebooted afterReb061015.png
    Reply  Thu Jun 11 19:40:59 2015, Koji, Update, VAC, Vacuum comp. rebooted 
       Reply  Fri Jun 12 08:40:17 2015, Steve, Update, VAC, Vacuum comp. rebooted vacMonNoGauges.pngprepReboot.png
Message ID: 11354     Entry time: Fri Jun 12 08:40:17 2015     In reply to: 11353
Author: Steve 
Type: Update 
Category: VAC 
Subject: Vacuum comp. rebooted 

Koji and Steve,

One computer expert and one vacuum expert required.

Quote:

The serial connections to the vacuum gauges were recovered by rebooting c1vac1 and c1vac2.

Steve claimed that the vacuum screen had showed "NO COMM" at the vacuum pressure values.
The epics connection to c1vac was fine. We could logged in to c1vac1 with telnet too although c1vac2 had no response.

After some inspection, we decided to reboot the slow machines. Steve manually XXXed YYY valves (to be described)
to prepare for any possible unwanted switching. Initially Koji thought only c1vac2 can be rebooted. But it was wrong.
If the reset button is pushed, all of the modules on the same crate is reset. So everything was reset. After ~3min we still
don't have the connection to c1vac1 restored. We decided to another reboot. This time I pushed c1vac1 reset button.
After waiting about two minutes, the ADCs started to show green lights and the switch box started scanning.
We recovered the telnet connection to c1vac1 and epics functions. c1vac2 is still note responding to telnet, and
the values associated with c1vac2 are still blank.

Steve restored the valves and everything was back to normal.

Atm 1,  problem condition: gauges are not reading for a week, error message "NO COMM" and all computer LEDs are green

Atm 2, prepare to safe reboot:

            a, close V1, disconnect it's power cable and turn off Maglev, wait till rotation stops

            b, close PSL shutter ( take adrenaline if needed )

            c, close V4, V5, VA6 valves and disconnect their cables. "Moving" error message indicating this condition.

               V1 is not showing "Moving" because its power cable disconnected only! It will show it if its position indicator cable is disconnected too. There is no need for that.

               These valves closed and disabled will not allow accidental venting of main volume.

            d, push reset, reseting c1vac2 will reset c1vac1 also, wait ~ 6 minutes

"Vacuum Normal" valve configuration was restored after succesful reboot as follows:

             a, reconnect cable and open V4 and V5 at P2 & P3 <1e-1 Torr

             b, observe that P2  < 1e-3 Torr and retsart Maglev

             c, wait till Maglev reaches full speed of 560 Hz and reconnect-open V1

             d, reconnect-open VA6 at P3 <1e-3 Torr

NOTE: VM1 valve was locked in open position and it was not responding before and after reboot

          Error message on Atm2 is indicating this locked condition: "opening VM1 will vent IFO"

          This is a fauls message. The valve is frozen in open position. We need a softwear expert help.

 

 

Attachment 1: vacMonNoGauges.png  34 kB  | Hide | Hide all
vacMonNoGauges.png
Attachment 2: prepReboot.png  34 kB  | Hide | Hide all
prepReboot.png
ELOG V3.1.3-